# Data-Driven Discovery of Digital Twins in Biomedical Research

### v1

@@ 4.2 Approaches for Handling Unobserved Derivatives
A fundamental challenge in inferring ODEs is the common unavailability of the vector field, as time derivatives are rarely experimentally measured. Consequently, methods fall into three broad categories: those that numerically estimate derivatives, those that approximate the candidate model solutions directly within the optimization procedure, and those that circumvent differentiation through reformulation.
@@
Alternatively, recent approaches numerically solve the candidate models into the inference process, enabling their direct comparison with state variable measurements, in an end-to-end learning. For example, He and Sun (2024) incorporate total variation regularization on the state trajectories during optimization. This encourages piecewise smooth solutions, which implicitly yield noise-robust derivatives without requiring their explicit computation. These regularized trajectories are then used internally for model fitting, tightly coupling trajectory estimation with parameter inference. More expressive frameworks rely on Gaussian Processes or neural ODEs to jointly learn both the latent dynamics and observational noise. A prominent example is GP-NODE (Bhouri and Perdikaris, 2022), which combines sparse identification of dynamics via SINDy, a differentiable neural ODE solver, and a Gaussian Process observation model. At each iteration, candidate dynamics (expressed as a sparse combination of nonlinear basis functions) are integrated via a neural ODE module to produce a continuous trajectory. This trajectory is then compared to the observed state variables time series and a Gaussian Process is used to model the mismatch: its mean is given by the neural ODE trajectory, and its covariance captures the residual uncertainty due to observation noise or model misspecification. These methods improve robustness but introduce challenges: higher computational cost, identifiability issues, and sensitivity to hyperparameters such as neural network depth or Gaussian process kernel choice.
The last class of methods avoids derivative estimation altogether by reformulating the inference problem. Integral formulations (Goyal and Benner, 2022; Anvari et al., 2025; Schaeffer and McCalla, 2017; Wu et al., 2025) rewrite the ODE as an integral equation, replacing differentiation with numerical integration. This makes the approach more robust to noise but introduces trade-offs such as increased computational cost and possible error accumulation over long trajectories. Weak formulations (Messenger and Bortz, 2021; Pantazis and Tsamardinos, 2019; Qian et al., 2022; Nicolaou et al., 2023; Fung et al., 2025) extend this idea by projecting the dynamics onto a family of smooth test functions. This projection further enhances stability and noise resilience, and extensive benchmarking found the weak integral formulation to consistently outperform other strategies, even in noiseless cases—though their evaluation was limited to chaotic systems (Kaptanoglu et al., 2023; Messenger and Bortz, 2021). However, design choices like selecting appropriate basis functions or integration intervals may significantly impact performance.
@@ 4.3 Selection, Refinement or Generation of Candidate Function Libraries
Candidate functions define the hypothesis space from which governing equations are inferred. In sparse regression, they form a predefined library of features; in symbolic regression, they emerge dynamically from combinations of variables, constants, and basic operators. While conceptually distinct, both approaches rely on suitable function sets to capture system dynamics. Overly limited candidate sets can hinder expressiveness, while overly large ones can overwhelm inference. To address this, various strategies select, refine, or adapt candidate functions using data-driven or knowledge-based heuristics. These fall into two categories: properly selecting the candidate function library before model learning, or refining it during model discovery.
@@
Methodological challenge 3 - Key takeaways Taken together, these strategies demonstrate a shift from static candidate sets to adaptive, data-informed function spaces that evolve alongside the model. Selecting or refining candidate functions prior to or during model learning may not only reduce the hypothesis space and accelerate computation, but also facilitate the discovery of more accurate and interpretable models (Naozuka et al., 2022). However, to our knowledge, no benchmark study has yet quantified the actual gains in terms of model quality, robustness, and computational efficiency provided by candidate library design. Yet challenges remain as many feature selection methods rely on ad hoc heuristics with limited theoretical grounding, and are based on a one-shot step rather than iterative workflows incorporating data fit residuals or expert feedback. While LLM-based approaches offer flexibility, they remain brittle, dependent on prompt quality, and are prone to producing overly complex or biologically implausible expressions (Sharlin and Josephson, 2024).
@@ 5.1 Strengths and limitations of existing symbolic regression methods in terms of identified challenges
Symbolic regression methods offer an interpretable framework for discovering compact analytical expressions, yet they struggle to meet the specific requirements of biological modeling as summarized in Table 1. Among the 49 best-performing symbolic regression methods reviewed here, only three of them address more than three challenges. Eleven out of 49 algorithms handle either "Missing data" or "Unobserved variables" challenges while only one method may incorporate both. In addition, none of the reviewed methods provides solutions for the challenge of integrating "Multiple experimental conditions". Similarly, symbolic regression methods largely do not meet key methodological criteria such as avoiding derivative estimation (8/49 methods) and uncertainty quantification (9/49 methods). Importantly, high-dimensional systems remain particularly challenging: only seven methods have been tested on systems with four or more state variables, and some studies report severe computational burdens in such settings (Cranmer, 2023; Yu et al., 2025a). These unresolved challenges are compounded by a structural limitation: most symbolic regression frameworks neglect explicit modeling of inter-variable dependencies, making most of them poorly suited for inferring coupled ODE systems in biology. Nonetheless, symbolic regression approaches show remarkable efforts in the generation and selection of candidate functions (22/49 methods), as detailed in Section 4.3. In parallel, 29 methods integrate prior knowledge, most often through pretrained neural networks or LLMs, offering complementary pathways for enhancing model discovery.
@@ 5.2 Strengths and limitations of existing sparse regression methods in terms of identified challenges
Among sparse regression approaches, the most effective methods that address the highest number of identified biological and methodological challenges were all Bayesian frameworks, demonstrating the superiority of such approach for model learning (Bhouri and Perdikaris (2022); Jiang et al. (2022); Long et al. (2024); Course and Nair (2023); Foo et al. (2025); Fung et al. (2025); North et al. (2022); Sun et al. (2022)). The methods by Bhouri and Perdikaris (2022), Jiang et al. (2022), and Long et al. (2024) emerge as the only pipelines addressing more than five out of the eight challenges considered in this review. All three methods i) embed interpolation or trajectory reconstruction techniques to mitigate the effects of data noise or missing values, ii) integrate prior knowledge through sparsity-enforcing parameter priors, iii) were validated on biological systems with at least 4 dimensions, iv) directly perform a numerical resolution of the ODE system to avoid the instability of finite differences-based derivative approximations, v) offer a quantification of model uncertainty through parameter posterior distribution. Bhouri and Perdikaris (2022) go further by handling unobserved variables (see section 4.2. To further improve these methods, several directions emerge from the review. For handling multiple experimental conditions, Bayesian sparse regression is compatible with multi-condition model discovery as in Schaeffer et al. (2017) and Park et al. (2023). Another solution, particularly when larger datasets are available, is to adopt a two-stage pipeline using the neural representation learning approach of Gui et al. (2025) to capture invariant functions and apply sparse Bayesian inference on those learned functions to extract interpretable mathematical structures. Finally, regarding candidate function selection, most documented sparse regression methods would benefit from incorporating statistical criteria upstream or during inference, as discussed in Section 4.3.
@@ 6 Discussion
In this review, we surveyed recent advances in the data-driven inference of biological dynamical systems represented as mechanistic ODEs, focusing on two main approaches —symbolic or sparse regression— summarized in an interactive Shiny application [^1] https://u1331systemspharmacology.shinyapps.io/model_learning_review/. We assessed how these methods address four key biological challenges: noisy and irregular data, heterogeneity across conditions, incorporation of prior biological knowledge, and partial observability of system states. We also considered four methodological issues associated with candidate function library design, scalability to high-dimensional systems, unobserved state variable derivatives, and uncertainty quantification. Overall, sparse regression emerges as more efficient than symbolic regression, particularly in Bayesian formulations, which excel at incorporating prior knowledge and quantifying uncertainty in both parameters and inferred models — crucial features in noisy, underdetermined biological settings. However, no inference method addresses more than six out of eight challenges, which advocates for future developments. Scalability to high-dimensional networks remains a main bottleneck, as candidate libraries grow combinatorially and data demands rise with model size.
Of utmost interest, deep learning methods are increasingly combined with sparse and symbolic regression, resulting in improved performance. In sparse regression, neural ODEs help denoise trajectories and capture latent dynamics before term selection (Bhouri and Perdikaris, 2022; Wu et al., 2025). In symbolic regression, hybrid neural ODEs blend neural approximators with domain constraints to guide expression search and improve interpretability (Grigorian et al., 2025; Bendinelli et al., 2023). These approaches complement foundation and large language models, which offer new ways to exploit prior knowledge from literature, structured databases, or expert input. As an important example, LLMs trained on ODE simulations can predict symbolic equations from incomplete trajectories, which can serve as an educated prior or to assign plausibility scores to candidate library terms based on biological context. However, their lack of explicit temporal modeling limits their ability to simulate dynamics, and challenges remain regarding their consistency, factual reliability, and bias, underscoring the need for rigorous evaluation and control.
@@ 7.1 Toward integrated and reliable frameworks
Overall, our analysis shows that no single model learning method currently addresses all the challenges of biological systems, with real-world applicability still limited. Since each challenge has been addressed in at least one study, the key insight is that progress is likely to come less from entirely new algorithms than from integrating existing ones into flexible, modular frameworks. To that end, we advocate for a CRN-based approach, which naturally encodes biological structure through stoichiometry, kinetics, and reaction directionality (Section 4.1). Effective frameworks should also prioritize reliability under adverse conditions—avoiding explicit differentiation, refining candidate functions, and rigorously quantifying uncertainty. Notably, few methods distinguish epistemic uncertainty arising from limited knowledge or model misspecification from aleatoric noise, limiting interpretability. Future work should assemble these components into general-purpose tools for dynamic model discovery in biology, a necessary step toward realizing the full potential of data-driven modeling in complex living systems.
Looking ahead, the most exciting opportunities lie in combining these approaches with emerging AI models. Promising directions include (i) hybrid pipelines that integrate mechanistic priors with LLM-guided candidate generation, (ii) retrieval-augmented strategies that anchor outputs in curated sources such as pathway databases, and (iii) tools that translate expert input or experimental context into executable constraints, making models more adaptive and user-driven.

### v2

@@ 4.2 Approaches for Handling Unobserved Derivatives
A fundamental challenge in inferring systems of ODEs is the common unavailability of the vector field, as time derivatives are rarely experimentally measured. Consequently, methods fall into three broad categories: those that numerically estimate derivatives, those that approximate the candidate model solutions directly within the optimization procedure, and those that circumvent differentiation through reformulation of the problem.
@@
Alternatively, recent approaches numerically solve the candidate models during the inference process, enabling their direct comparison with state variable measurements, in an end-to-end pipeline. For example, He and Sun (2024) incorporate total variation regularization on the state trajectories during optimization. This encourages piecewise smooth solutions, which implicitly yield noise-robust derivatives without requiring their explicit computation. These regularized trajectories are then used internally for model fitting, tightly coupling trajectory estimation with parameter inference. More expressive frameworks rely on Gaussian Processes or neural ODEs to jointly learn both the latent dynamics and observational noise. A prominent example is GP-NODE (Bhouri and Perdikaris, 2022), which combines sparse identification of dynamics via SINDy, a differentiable neural ODE solver, and a Gaussian Process observation model. At each iteration, candidate dynamics (expressed as a sparse combination of nonlinear basis functions) are integrated via a neural ODE module to produce a continuous trajectory. This trajectory is then compared to the observed state variables time series and a Gaussian Process is used to model the mismatch: its mean is given by the neural ODE trajectory, and its covariance captures the residual uncertainty due to observation noise or model misspecification. These methods improve robustness but introduce challenges: higher computational cost, identifiability issues, and sensitivity to hyperparameters such as neural network depth or Gaussian process kernel choice.
The last class of methods avoids derivative estimation by reformulating the inference problem. In integral formulations (Goyal and Benner, 2022; Anvari et al., 2025; Schaeffer and McCalla, 2017; Wu et al., 2025), the ODE is rewritten as an integral equation, replacing differentiation with numerical integration. This makes the approach more robust to noise but introduces trade-offs such as increased computational cost and possible error accumulation over long trajectories. Weak formulations (Messenger and Bortz, 2021; Pantazis and Tsamardinos, 2019; Qian et al., 2022; Nicolaou et al., 2023; Fung et al., 2025) extend this idea by projecting the dynamics onto a family of smooth test functions. This projection further enhances stability and noise resilience, and extensive benchmarking found the weak integral formulation to consistently outperform other strategies, even in noiseless cases—though their evaluation was limited to chaotic systems (Kaptanoglu et al., 2023; Messenger and Bortz, 2021). However, design choices like selecting appropriate basis functions or integration intervals may significantly impact performance.
@@ 4.3 Selection, Refinement or Generation of Candidate Function Libraries
Candidate functions define the hypothesis space from which governing equations are inferred. In sparse regression, they form a predefined library of features; in symbolic regression, they emerge dynamically from combinations of variables, constants, and basic operators. While conceptually distinct, both approaches rely on suitable function sets to capture system dynamics. Overly limited candidate sets can hinder expressiveness, while overly large ones can overwhelm inference. To address this, various strategies offer to select, refine, or adapt candidate functions using data-driven or knowledge-based heuristics. These fall into two categories: properly selecting the candidate function library before model learning, or refining it during model discovery.
@@
Methodological challenge 3 - Key takeaways Taken together, these strategies demonstrate a shift from static candidate sets to adaptive, data-informed function spaces that evolve alongside the model inference. Selecting or refining candidate functions prior to or during model learning may not only reduce the hypothesis space and accelerate computation, but also facilitate the discovery of more accurate and interpretable models (Naozuka et al., 2022). However, to our knowledge, no benchmark study has quantified the actual gains in terms of model quality, robustness, and computational efficiency yet. Challenges remain as many feature selection methods rely on ad hoc heuristics with limited theoretical grounding, and are based on a one-shot step rather than iterative workflows incorporating data fit residuals or expert feedback. While LLM-based approaches offer flexibility, they remain brittle, dependent on prompt quality, and are prone to producing overly complex or biologically implausible expressions (Sharlin and Josephson, 2024).
@@ 5.1 Strengths and limitations of existing symbolic regression methods in terms of identified challenges
Symbolic regression methods offer an interpretable framework for discovering compact analytical expressions, yet they struggle to meet the specific requirements of biological modeling as summarized in Table 1. Among the 49 best-performing symbolic regression methods reviewed here, only three of them address more than three challenges out of the eight identified ones. Eleven out of 49 algorithms handle either "Missing data" or "Unobserved variables" while only one method may incorporate both. In addition, none of the reviewed methods provides solutions for the challenge of integrating "Multiple experimental conditions". Similarly, symbolic regression methods largely do not meet key methodological criteria such as avoiding derivative estimation (8/49 methods) and uncertainty quantification (9/49 methods). Importantly, high-dimensional systems remain particularly challenging: only seven methods have been tested on systems with four or more state variables, and some studies report severe computational burdens in such settings (Cranmer, 2023; Yu et al., 2025a). These unresolved challenges are compounded by a structural limitation: most symbolic regression frameworks neglect explicit modeling of inter-variable dependencies, making most of them poorly suited for inferring coupled ODE systems. Nonetheless, symbolic regression approaches show remarkable efforts in the generation and selection of candidate functions (22/49 methods), as detailed in Section 4.3. In parallel, 29 methods integrate prior knowledge, most often through pretrained neural networks or LLMs, offering complementary pathways for enhancing model discovery.
@@ 5.2 Strengths and limitations of existing sparse regression methods in terms of identified challenges
Among sparse regression approaches, the most effective methods that address the highest number of identified biological and methodological challenges were all Bayesian frameworks, demonstrating the superiority of such approach for model learning (Bhouri and Perdikaris (2022); Jiang et al. (2022); Long et al. (2024); Course and Nair (2023); Foo et al. (2025); Fung et al. (2025); North et al. (2022); Sun et al. (2022)). The methods by Bhouri and Perdikaris (2022), Jiang et al. (2022), and Long et al. (2024) emerge as the only pipelines addressing more than five out of the eight challenges considered in this review. All three methods i) embed interpolation or trajectory reconstruction techniques to mitigate the effects of data noise or missing values, ii) integrate prior knowledge through sparsity-enforcing parameter priors, iii) were validated on biological systems with at least 4 dimensions, iv) directly perform a numerical resolution of the ODE system to avoid the instability of finite differences-based derivative approximations, v) offer a quantification of model uncertainty through parameter posterior distribution. Bhouri and Perdikaris (2022) go further by handling unobserved variables (see section 4.2). To further improve these methods, several directions emerge from the review. For handling multiple experimental conditions, Bayesian sparse regression is compatible with multi-condition model discovery as in Schaeffer et al. (2017) and Park et al. (2023). Another solution, particularly when larger datasets are available, is to adopt a two-stage pipeline using the neural representation learning approach of Gui et al. (2025) to capture invariant functions and apply sparse Bayesian inference on those learned functions to extract interpretable mathematical structures. Finally, regarding candidate function selection, most documented sparse regression methods would benefit from incorporating statistical criteria upstream or during inference, as discussed in Section 4.3.
@@ 6 Discussion
In this review, we surveyed recent advances in the data-driven inference of biological digital twins represented as mechanistic ODEs, focusing on two main approaches —symbolic or sparse regression— summarized in an interactive Shiny application [^1] https://u1331systemspharmacology.shinyapps.io/model_learning_review/. We assessed how these methods address four key biological challenges: noisy and irregular data, heterogeneity across conditions, incorporation of prior biological knowledge, and partial observability of system states. We also considered four methodological issues associated with scalability to high-dimensional systems, unobserved state variable derivatives, candidate function library design, and model uncertainty quantification. Overall, sparse regression emerges as more efficient than symbolic regression, particularly in Bayesian formulations, which excel at incorporating prior knowledge and quantifying uncertainty in both parameters and inferred models — crucial features in noisy, underdetermined biological settings. However, no inference method addresses more than six out of the eight challenges, which advocates for future developments. Scalability to high-dimensional networks remains a main bottleneck, as candidate libraries grow combinatorially and data demands rise with model size.
Of utmost interest, deep learning methods are increasingly combined with sparse and symbolic regression, resulting in improved performance. In sparse regression, neural ODEs help denoise trajectories and capture latent dynamics before term selection (Bhouri and Perdikaris, 2022; Wu et al., 2025). In symbolic regression, hybrid neural ODEs blend neural approximators with domain constraints to guide expression search and improve interpretability (Grigorian et al., 2025; Bendinelli et al., 2023). These approaches complement foundation and large language models, which offer new ways to exploit prior knowledge from literature, structured databases, or expert input. As an important example, LLMs trained on ODE simulations can predict symbolic equations from incomplete trajectories, which can serve as an educated prior or for assigning plausibility scores to candidate library terms based on biological context. However, their lack of explicit temporal modeling limits their ability to simulate dynamics, and challenges remain regarding their consistency, factual reliability, and bias, underscoring the need for rigorous evaluation and control.
@@ 7.1 Toward integrated and reliable frameworks
Overall, our analysis shows that no single model learning method currently addresses all the challenges of biological systems, with real-world applicability being still limited. Since each challenge has been addressed in at least one study, the key insight is that progress is likely to come less from entirely new algorithms than from integrating existing ones into flexible, modular frameworks. To that end, we advocate for a CRN-based approach, which naturally encodes biological structure through stoichiometry, kinetics, and reaction directionality (Section 4.1). Effective frameworks should also prioritize reliability under adverse conditions—avoiding explicit differentiation, refining candidate functions, and rigorously quantifying model uncertainty. Notably, few methods currently distinguish epistemic uncertainty arising from limited knowledge or model misspecification from aleatoric noise, limiting interpretability. Future work should assemble these components into general-purpose tools for dynamic model discovery in biology, a necessary step toward realizing the full potential of data-driven modeling in complex living systems.
Looking ahead, the most exciting opportunities lie in combining these approaches with emerging AI models. In this area, promising directions include (i) hybrid pipelines that integrate mechanistic priors with LLM-guided candidate generation, (ii) retrieval-augmented strategies that anchor outputs in curated sources such as pathway databases, and (iii) tools that translate expert input or experimental context into executable constraints, making models more adaptive and user-driven.
@@ 8 Methods
9 Acknowledgment
We sincerely thank Lucie Gaspard-Boulinc (Institut Curie, INSERM U1331) for her contribution in the development of the Shiny application and for her help with the visualization of the results of this review.